Determine whether the graphs of the pair of lines are parallel.
x+8=y
y-x=-4
What is the slope of the line x+8=y?
What is the slope of the line y-x=-4?
Are the graphs of the given lines parallel?



The function p(d)=1+d/33 gives the pressure, in atmospheres (atm), at a depth d in the sea (d is in feet). Note that p(0) = 1atm, p(33)=2 and so on. Find the pressure at 150 feet.

Answer by Alan3354(69443)

Determine whether the graphs of the pair of lines are parallel.
x+8=y
y-x=-4
What is the slope of the line x+8=y?
y = 1x + 8 What is the slope?
--------------------
What is the slope of the line y-x=-4?
y = 1x - 4 What is the slope?
---------------------------
Are the graphs of the given lines parallel?
If the slopes are the same, they're parallel.
-------------------
The function p(d)=1+d/33 gives the pressure, in atmospheres (atm), at a depth d in the sea (d is in feet). Note that p(0) = 1atm, p(33)=2 and so on. Find the pressure at 150 feet.
p(150) = 1 + 150/33